( 1 ) ( 2 ) ( 3 ) ( 4 ) ( 5 )
What are the odds in FAVOR of drawing a 2 from these numbers?
I have 1 out of 5 Is that right??

Answer by stanbon(48568)

Odds in favor of A = P(A)/P(A')
P(2)=1/5
P(2')=4/5
Odds in favor of "2" = 1:4
Cheers,
Stan H.